NASA ISS On-Orbit Status 24 February 2015


NASA ISS On-Orbit Status 24 February 2015.

Today: Extravehicular Activity (EVA) #30 Preparation: The USOS crew members made final preparations for the tomorrow's EVA. Today, they reviewed the detailed EVA timeline and translation paths, completed an EVA tool audit, and configured the equipment lock.

Tomorrow's EVA will complete the IDA-1 cable installation and accomplish Space Station Remote Manipulator System (SSRMS) Latching End Effector (LEE) Lubrication. Joint Airlock Egress is scheduled to occur at ~6:15am CST with ingress taking place at ~12:10pm CST.

Urine Processing Assembly (UPA) Status: Flight Controllers were not able to perform a process cycle yesterday evening due to high pressure within the system, following the Fluids Control Pump Assembly (FCPA) change out. Today, ground teams worked to lower the system pressure to the required 24 mmHg. Once the targeted pressure was reached, a process cycle was attempted. However, UPA faulted out shortly afterwards. The Waste and Hygiene Compartment (WHC) remains configured for internal tank (EDV), while ground teams investigate the latest UPA error code.

External (EXT)-2 Multiplexer/Demultiplexer (MDM): This morning the primary EXT MDM (EXT-2) lost communication with the upstream Primary Command and Control (C&C) MDM. Redundancy Management Software executed as expected and removed power from the MDM then commanded EXT-1 MDM as primary. Since then, Flight Controllers have successfully re-powered EXT-2 MDM and are receiving good health and status from the MDM with no failures being reported. Ground teams will continue to monitor EXT-2 and investigate the loss of communication event. The EXT MDM is responsible for specific control and data processing of multiple external systems, including the Mobile Transporter (MT), Passive Thermal Control System (PTCS), Solar Alpha Rotary Joint (SARJ), Thermal Radiator Rotary Joint (TRRJ), and the Secondary Electrical Power System (SEPS).

Potable Water Dispenser (PWD) Filter ORU Change Out: As part of periodic maintenance, Cristoforetti changed out the PWD Filter today.

Today's Planned Activities
All activities are completed unless otherwise noted.

OBSTANOVKA. Download of Obstanovka P/L Data from БСПН (Payload Server) to RSS1 Laptop
CASKAD. Manual Mixing in Bioreactor
Start Battery Maintenance
EHIP - Hardware Transfer to RS crew to be used for FGB СЭС and СУБК Units Ops
D2XS Battery Charge
Installation of REBA batteries
TOCA Waste Water Bag (WWB) Changeout
PROBOY. Temporary Transfer of РМ1 Microphone to SM [РО]
WRS Water Sample Analysis
REBA powered equipment test
Pre-EVA Crew Health Status - Prep
PROBOY. RSЕ1 Station Laptop and Otklik Hardware Ops
PROBOY. Penetration Simulator Ops
Pre-EVA Crew Health Status [Aborted per Surgeon and crew's discretion]
SM CO measurement using CMS unit
PROBOY. Installation of РМ1 Microphone
Pre-EVA Crew Health Status Data Entry and Stow
OTKLIK. Hardware Check
EMU Metal Oxide (METOX) Canister Installation
Scheduled Maintenance of Compound Specific Analyzer- Combustion Products (CSA-CP)
OBSTANOVKA. Download of Obstanovka P/L Data from БСПН (Payload Server) to RSS1 Laptop (terminate) / r/g 7669
MATRYOSHKA-R. Installation of Personal Hygiene Articles СЛГ and СЛГ-3 into starboard CQ protective curtain pockets and installation of the curtain
USOS EVA Tool Configuration
Conference between Search and Rescue (ГПСК) specialists and the returning crew.
Potable Water Dispenser (PWD): Filter Removal and Replacement (R&R)
Troubleshooting FGB СУБК Instruments 11М156М
VIBROLAB. Experiment Timeline Review
TOCA Data Recording
VIBROLAB. Re-mating connectors, photography
VIBROLAB. SINUS-ACCORD P/L Activation, Mode 1 Activation
Testing FGB СЭС БФ-2 and БСШ-2
VIBROLAB. Switching from Mode-1 to Mode-2
D2XS Battery Charge
DOUG Software Review
On-orbit Hearing Assessment using EARQ
PILLE Data Download & placement before EVA
Data Transmission Radio Link (РСПИ) Installation of SW for РСПИ Log files auto prep and downlink
USOS EVA Tool Audit
VIBROLAB. SINUS-ACCORD P/L Deactivation. Data downlink
USOS EVA Procedure Review
VIBROLAB. Telemetry Downlink Test
Testing FGB СЭС БФ-2 and БСШ-2
USOS EVA Procedure Conference
SPLANH. Experiment Timeline Review
Equipment Lock Preparation
On-orbit Hearing Assessment using EARQ
DOUG Software Review
IMMUNO. Preparing Saliva-Immuno Kit for the experiment
DOUG Conference
Medical Ops OBT
D2Xs EVA Camera Configuration for EVA
IMMUNO. Preparing Saliva-Immuno Kit for the experiment
CASKAD. Manual Mixing in Bioreactor
ABOUT GAGARIN FROM SPACE. Equipment deactivation and cleanup
Completed Task List Items

Ground Activities
All activities are completed unless otherwise noted.

ELC3 software update
Three-Day Look Ahead:

Wednesday, 02/25: EVA #30 IDA Cable / Lee Lube
Thursday, 02/26: EVA Debrief, EVA Tool Configuration, JEMAL depressurization
Friday, 02/27: Microbe Sample, Triplelux, NRCSD Deployment, EVA Tool Config, Plant Rotation, OBT 41S Emer Drill, WHC Pre-treat tank R&R
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - On
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Idle
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Standby
Urine Processing Assembly (UPA) - Shutdown
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Please follow SpaceRef on Twitter and Like us on Facebook.